OpenStudy (anonymous):

Given f(x)=1-x+sinx ; Approximate the zero(s) of the function. Use Newton’s Method where x1=2 and continue the process until two successive approximations differ by less than 0.001.

OpenStudy (anonymous):

Sorry not to actually be answering your question, but dang, it looks like you stole my cat! XD

ganeshie8 (ganeshie8):

\[x_{n+1} = x_n - \dfrac{f(x_n)}{f'(x_n)}\]

OpenStudy (anonymous):

@ganeshie8 ok so whats next? is n equal to 0.001?

ganeshie8 (ganeshie8):

your first guess is \(x_1 = 2\)

ganeshie8 (ganeshie8):

newton's method gives u the next good guess : \[x_{1+1} = x_1 - \dfrac{f(x_1)}{f'(x_1)} \]

ganeshie8 (ganeshie8):

evaluate \(f(x_1)\) and \(f'(x_1)\) , plug them above ^

OpenStudy (anonymous):

i have to go but ill be on tomorrow to review :( thanks!

ganeshie8 (ganeshie8):

okay sure :)